### v1

# Optimal Transportation by Orthogonal Coupling Dynamics

Abstract.
Many numerical algorithms and learning tasks rest on solution of the Monge-Kantorovich problem and corresponding Wasserstein distances. While the natural approach is to treat the problem as an infinite-dimensional linear programming, such a methodology severely limits the computational performance due to the polynomial scaling with respect to the sample size along with intensive memory requirements. We propose a novel alternative framework to address the Monge-Kantorovich problem based on a projection type gradient descent scheme. The micro-dynamics is built on the notion of the conditional expectation, where the connection with the opinion dynamics is explored and leveraged to build compact numerical schemes. We demonstrate that the devised dynamics recovers random maps with favourable computational performance. Along with the theoretical insight, the provided dynamics paves the way for innovative approaches to construct numerical schemes for computing optimal transport maps as well as Wasserstein distances.
Corresponding author: Hossein Gorji.
Emails: mohsen.sadr@psi.ch, P.MohajerinEsfahani@tudelft.nl, Mohammadhossein.Gorji@empa.ch. Mohsen Sadr: Paul Scherrer Institute, Forschungsstrasse 111, CH-5232 Villigen, Switzerland. Peyman Mohajerin Esfehani: The Delft Center for Systems and Control, Delft University of Technology, Delft, The Netherlands. Hossein Gorji: Laboratory for Computational Engineering, Empa, Dübendorf, Switzerland.
1. Introduction
Background. The importance of the optimal transport lies in its widespread application at different fronts of computational studies, along with its unique theoretical properties. It plays an essential role in machine-learning by offering relevant metrics for comparing different distributions [5, 39, 26]. It has led to a significant progress in the theory of partial differential equations by its by-product, Wasserstein gradient flows [2, 35, 57]. At the same time, existing algorithms to compute such distances and corresponding maps remain expensive and complex. Given the versatile role of optimal transport in different branches of machine-learning [34, 1, 5], analysis [63, 35, 56], density functional theory [18, 24], optimization [43], and inference [32], among others, several numerical approaches have been pursued for its efficient computations. Classical methods were developed based on linear programming [13], gradient descent [19], dynamic flow formulation [10, 62], or elliptic solvers [58]. For efficient computation and improved performance-scaling with respect to the dimension, entropy regularization [25, 51] and its stochastic treatment [7] have been pursued. Approximate approaches using moment formulation [45, 55] of Monge-Kantorovich problem have been proposed, as well.
Despite these progresses, to the best of our knowledge, there is no concise Ordinary Differential Equation (ODE) model, in the sense of particle/agent based models, which leads to the solution of the (un-regularized) Monge-Kantorovich problem. Such a model would be particularly valuable, beyond theoretical and conceptual contributions, for its potential to further enable efficient optimal transport computations. In the opening chapter of Villani’s seminal book [63], a list of coupling dynamics, of ODE or Stochastic Differential Equation (SDE) forms, are reviewed with the intention of creating correlations between two given random variables. While interesting results can be obtained e.g. by the Knoth-Rosenblatt rearrangement [28] or Moser’s coupling [17], such schemes remain limited; e.g. due to coordinate dependency of the Knoth-Rosenblatt rearrangement, or the constraint of small deviation between the marginals for Moser’s coupling.
In principle, finding the optimal transport between two probability spaces is a global optimization in a sense of linear programming on an infinite dimensional space. We are confronted with the question whether there exist local dynamic rules that can lead to the optimal transportation between two sample spaces? An interesting class of dynamic processes is given by orthogonal dynamics. They naturally arise in projecting Hamiltonian systems onto a sub-space, e.g. by the use of conditional expectation [21, 31]. These processes have an interesting feature that the remainder/unresolved portion remains orthogonal to the current state of the resolved variable.
Recent study of Conforti et al. [23] leverages the projection method in the context of Langevin dynamics. Their proposed stochastic dynamics leads to the coupling which converges to the Sinkhorn regularization of the optimal transport problem [25]. Such a stochastic dynamics, which is closely linked to the Schrödinger Bridge [30, 40, 20], gives rise to the Fokker-Planck equation, evolving the joint distribution. It remains to be addressed what happens to the introduced projected Langevin dynamics when the contribution of Brownian motion vanishes, relevant for the (un-regularized) Monge-Kantorovich problem.
The simplest form of the orthogonal dynamics can be constructed by projection of the gradient descent (with respect to the Monge-Kantorovich cost) onto a sub-space where marginals are preserved. In the absence of entropy regularization, this dynamics becomes a good candidate to create optimal transport between two measure spaces. At the level of distribution, one expects such a dynamics gives rise to the Vlasov equation for the evolution of the joint. Despite the theoretical appeal of the coupling through projection, numerical treatment of the conditional expectation can become prohibitive. In general, conditional expectation is solution of the Bayesian regression [46]. On the other hand, interpreting the conditional expectation through distinct data clusters, e.g. in an analogy with opinion dynamics [41], offers non-parametric numerical approach to deal with such coupling dynamics. With these observations in hand and leveraging the recent study [23], we pursue a projected dynamic approach as a concise solution algorithm of the Monge-Kantorovich problem.
Main Contributions. We present the orthogonal coupling dynamics for the Monge-Kantorovich problem. More specifically:
  1. (1)
    Main Dynamics. In Sec. 2.1, we introduce the dynamics in Eq. (2.1), motivated by projection of the gradient descent on a marginal-preserving tangent space.
  2. (2)
    General Structural Properties. In Sec. 2.2, assuming that the well-posedness of the dynamics is granted, we justify several structural properties for the general cost setting.
    1. (a)
      Marginal Preservation. We prove in Proposition 7.1 that the introduced dynamics preserves the marginals.
    2. (b)
      Descent in Cost. The dynamics has a monotone behaviour in decreasing the cost, as shown in Proposition 2.2.
    3. (c)
      Instability of Sub-Optimal Couplings. We find that the sub-optimal maps are unstable as proved in Proposition 2.3.
    4. (d)
      McKean-Vlasov System. We derive the corresponding kinetic description of the introduced evolution in Theorem 2.4, where the kinetics takes the form of the Vlasov equation (20).
  3. (3)
    Key Features in $L^{2}$-Monge-Kantorovich. Next, we turn our attention to the common case of $L^{2}$-cost and derive two results in Sec. 2.3.
    1. (a)
      Symmetric Positive Semi-Definite Correlation. We observe that the dynamics keeps the centered cross-correlation between samples of the two measure spaces in the cone of symmetric positive semi-definite matrices, as proved in Proposition 2.5.
    2. (b)
      Sharp Descent and Variational Formulation. We derive the variational form of the dynamics, which according to Theorem 2.6 gives the sharpest descent in the $L^{2}$-Monge-Kantorovich cost among a class of marginal-preserving dynamics.
  4. (4)
    Special Examples. In Sec. 3, two special cases of linear maps and elliptic distributions, where explicit results can be obtained, are discussed.
  5. (5)
    Numerical Algorithm. In Sec. 4, after motivating the analogy with opinion dynamics, a parameter-free recipe is proposed for computing the conditional expectation.
  6. (6)
    Numerical Experiments. We perform several numerical experiments to verify the convergence of the dynamics, as well as its relevance in practical settings, in Sec. 5. We demonstrate that the presented dynamics, equipped with the devised numerical algorithm, successfully recover nonlinear Monge maps. Furthermore, we showcase its performance in distribution learning, data-set classification, and color interpolation.
Our focus remains on basic structural properties of the dynamics along with its computational implications. The aim of the study therefore should be understood as demonstrating the relevance of the orthogonal coupling dynamics, and what it entails, rather than providing a rigorous comprehensive theory, or a thorough numerical analysis, of the presented dynamics itself.
Notation and Setup. Let $\mathcal{P}_{2}(\mathbb{R}^{n})$ be the set of Borel measures over $\mathbb{R}^{n}$ with finite second moment, i.e. $\mathcal{P}_{2}(\mathbb{R}^{n})\mathrel{\mathop{:}}=\{\mu\in\mathcal{P}_{2}(\mathbb{R}^{n})|\int_{\mathbb{R}^{n}}\lVert x\rVert_{2}^{2}\mu(dx)<\infty\}$, where .$\lVert.\rVert_{2}^{2}$ is the usual $L^{2}$ Euclidean norm. Suppose $\Pi(\mu,\nu)$ is the set of all joint measures in $\mathcal{P}_{2}(\mathbb{R}^{n}\times\mathbb{R}^{n})$ with marginals $\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{n})$. We denote by $\pi^{\textrm{opt}}$ the solution of the Monge-Kantorovich problem
[…]
with $c:\mathbb{R}^{2n}\to\mathbb{R}$ as a non-negative smooth function (see e.g. [12] for details). For the standard setting of $c(x,y)=\lVert x-y\rVert_{2}^{2}$, the optimal map $\hat{T}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ exists, if $\mu$ and $\nu$ are absolutely continuous with respect to the Lebesgue measure [2]. It gives $\hat{T}\#\mu=\nu$ (##\## is the push-forward) and accordingly the 2-Wasserstein distance reads
$d^{2}(\mu,\nu)$ $=$ $\int_{\mathbb{R}^{n}\times\mathbb{R}^{n}}\lVert x-y\rVert_{2}^{2}\pi^{\textrm{opt}}_{\mu,\nu}(dx,dy)$ (2)
$=$ $\int_{\mathbb{R}^{n}}\lVert\hat{T}(x)-x\rVert_{2}^{2}\mu(dx)\ .$ (3)
Let $\mathcal{H}=\mathcal{L}^{2}(\Omega,\mathscr{A},P)$ be the Hilbert space comprised of square-integrable $\mathscr{A}$-measurable functions on $\Omega$. Suppose $X_{t},Y_{t}:\Omega\to\mathbb{R}^{n}$ are random processes in $\mathcal{H}$ with the index $t\in[0,+\infty)$ and joint law $p_{t}$. We employ the conditional expectation $\mathbb{E}_{p_{t}}[X_{t}|Y_{t}]$ as the orthogonal projection of $X_{t}$ onto the span of $\sigma(Y_{t})$-measurable functions in $\mathcal{H}$, denoted by $\mathsf{P}_{{p_{t}},{Y_{t}}}[X_{t}]$, where $\sigma(A)$ is the smallest $\sigma$-algebra with respect to which the random variable $A$ is measurable (see e.g. [22, 14] for further details). Accordingly, we have
$\mathsf{P}_{p_{t},X_{t}}[Y_{t}]=\mathbb{E}_{p_{t}}[Y_{t}|X_{t}]\ \ \ \textrm{and}\ \ \ \mathsf{P}_{p_{t},Y_{t}}[X_{t}]=\mathbb{E}_{p_{t}}[X_{t}|Y_{t}]\ .$ (4)
We consider the probability flows following the continuity equation of the form
$\partial_{t}p_{t}+\nabla\cdot(vp_{t})$ $=$ $0\ ,$ (5)
initialized by $p_{0}\in\Pi(\mu,\nu)$, where the velocity vector $v=(v_{1},v_{2})$ is, in general, a functional of $p_{t}$ and belongs to the tangent space [3]
[…]
and therefore the tangent space is a Hilbert space comprised of gradient vector fields. Besides $\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$, we specifically consider two tangent spaces $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$ and $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$. The former, following [23], reads
$\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)\mathrel{\mathop{:}}=\left\{v\in\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n}):\int_{\mathbb{R}^{2n}}v\cdot(\nabla h_{1},\nabla h_{2})^{T}p_{t}(dx,dy)=0,\forall h_{1},h_{2}\in C_{c}^{\infty}(\mathbb{R}^{n})\right\}\ ,$ (7)
which gives the set of all velocities that preserve the marginals $\mu$ and $\nu$. Following notation of [23], the latter tangent space is given by
[…]
which is a sub-manifold of $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$, and as shown in Proposition 2.1, it preserves the marginals $\mu$ and $\nu$. Since for $v\in\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$, we have $\mathbb{E}_{p_{t}}[v_{1}(X_{t},Y_{t})|X_{t}]=\mathbb{E}_{p_{t}}[v_{2}(X_{t},Y_{t})|Y_{t}]=0$, the projection operator $\mathsf{S}_{p_{t}}:\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})\to\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ is orthogonal to $\mathsf{P}_{p_{t}}$, i.e.
$\mathsf{S}_{p_{t},X_{t}}[v_{1}(X_{t},Y_{t})]$ $=$ $v_{1}(X_{t},Y_{t})-\mathbb{E}_{p_{t}}[v_{1}(X_{t},Y_{t})|X_{t}]$ (9)
$\textrm{and}\ \ \ \mathsf{S}_{p_{t},Y_{t}}[v_{2}(X_{t},Y_{t})]$ $=$ $v_{2}(X_{t},Y_{t})-\mathbb{E}_{p_{t}}[v_{2}(X_{t},Y_{t})|Y_{t}]\ ,$ (10)
for $v\in\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$.
2. Orthogonal Coupling
2.1. Main Dynamics
Starting with independent samples of $X_{0}\sim\mu$ and $Y_{0}\sim\nu$, we would like to construct a micro-dynamics in the sense of an ODE on $X_{t}$ and $Y_{t}$, which converges to the solution of (1). In other words, we look for a velocity vector $v$ which updates $X_{t}$ and $Y_{t}$ in a way that the minimum cost $c(X_{t},Y_{t})$, in the average sense, is attained as $t\to\infty$. One can consider the conventional gradient descent for $X_{t}$ and $Y_{t}$ with respect to $c$, suggesting $v_{1}^{\textrm{gd}}=-\nabla_{x}c$ and $v_{2}^{\textrm{gd}}=-\nabla_{y}c$, and updating $X_{t}$ and $Y_{t}$ according to $v_{1}^{\textrm{gd}}$ and $v_{2}^{\textrm{gd}}$, respectively. However, it is easy to spot that the distributions of $X_{t}$ and $Y_{t}$ would go out of the set of couplings $\Pi(\mu,\nu)$.
An interesting approach to keep the marginals in $\Pi(\mu,\nu)$ is to project the dynamics of the gradient descent onto the tangent space which preserves the marginals. In particular, if we consider the probability flow given by
$\partial_{t}p_{t}+\nabla\cdot(\hat{v}p_{t})$ $=$ $0\ ,$ (11)
and restrict the velocities to those $\hat{v}=(\hat{v}_{1},\hat{v}_{2})\in\textrm{Tan}_{p_{t}}\mathcal{P}_{2}(\mathbb{R}^{2n})$ which fulfill
$\int_{\mathbb{R}^{2n}}\left(\hat{v}_{1}\cdot\nabla h_{1}(x)+\hat{v}_{2}\cdot\nabla h_{2}(y)\right)p_{t}(dx,dy)$ $=$ $0\,\ \ \ \forall h_{1,2}\in C^{\infty}(\mathbb{R}^{n}),$ (12)
the marginals of $p_{t}$ remain unchanged. However, the general form of this tangent space, given by $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$ in Eq. (7), does not provide us with a concise and convenient projection (this avenue was explored, to some degree, in [4]). Yet, if instead of $\textrm{Tan}_{p_{t},1}\Pi(\mu,\nu)$, we focus on a sub-manifold, which still preserves the marginals, the problem can become tractable. In particular, the tangent space $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)=(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}$ offers a straight-forward projection, while preserving the marginals. Noting that the Riemannian gradient is the orthogonal projection of the Euclidean one onto the tangent space [15], we get the projected version of $v^{\textrm{gd}}$, as shown in Fig. 1.
Figure 1. Geometric illustration of OCD; (a) The gradient descent, which leaves the polytope $\Pi(\mu,\nu)$, is projected onto the tangent space $(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}$; (b) Initial condition as product Gaussian probability density; (c) Evolved by $v^{\textrm{ocd}}$, the initial density is updated by projected gradient descent in $L^{2}$-cost; (d) Evolved by $v^{\textrm{gd}}$, the initial density is updated by gradient descent in $L^{2}$-cost, leaving $\Pi(\mu,\nu)$.
In fact, projecting $v^{gd}$ onto $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ gives us the update rule
$\dot{X}_{t}=v_{1}^{\textrm{ocd}}\mathrel{\mathop{:}}=-\nabla_{x}c(X_{t},Y_{t})+\mathbb{E}_{p_{t}}[\nabla_{x}c(X_{t},Y_{t})|X_{t}]$
$\textrm{and}\ \ \ \ \dot{Y}_{t}=v_{2}^{\textrm{ocd}}\mathrel{\mathop{:}}=-\nabla_{y}c(X_{t},Y_{t})+\mathbb{E}_{p_{t}}[\nabla_{y}c(X_{t},Y_{t})|Y_{t}]\,$ (13)
which we refer to as the Orthogonal Coupling Dynamics (OCD). The orthogonality comes from the observation that $\mathbb{E}[v_{1}^{\textrm{ocd}}\cdot X_{t}]=\mathbb{E}[v_{2}^{\textrm{ocd}}\cdot Y_{t}]=0$. In general, there is no guarantee that OCD reaches global minimum of (1), as the projection onto the sub-manifold $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ can prevent the dynamics from visiting all possible couplings between $\mu$ and $\nu$. This can be readily noticed, e.g. if $Y_{0}$ is initialized as a sub-optimal map of $X_{0}$. As a result, the dynamics remains frozen at a sub-optimal fixed point. However, as we will discuss, there are several nuances in (2.1), which warrant investigation of OCD on both theoretical and numerical fronts. In particular, we observe that OCD ensures the cost is non-increasing. This result plays a crucial role in two important properties that we deduce: we observe that sub-optimal couplings turn out to be unstable (see Proposition 2.3), and that OCD gives the sharpest decay in $L^{2}$-cost among processes with velocity in $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$ (see Theorem 2.6). In practice, the dynamics reaches accurate estimations of the optimization problem (1), as we demonstrate in results section.
From conceptual point of view, by replacing the optimization problem (1) with the OCD system (2.1), we turn the Monge-Kantorovich problem into a series of regression problems, implied by the conditional expectation. While the original problem, given by Eq. (1), is a constrained optimization over distributions, the OCD entails an unconstrained optimization over functions, in $L^{2}$. This transformation allows us to leverage tools of conditional expectation estimation in order to achieve efficient computation of optimal transport in contrast to the original constrained infinite dimensional linear programming problem.
However, before analyzing OCD, let us mention that alternatively, one can consider the Langevin processes
$d{X}_{t}$ $=$ $\left(-\nabla_{x}c+\mathbb{E}_{p_{t}}[\nabla_{x}c|X_{t}]+\epsilon\nabla_{x}\log f_{\mu}(X_{t})\right)dt+\sqrt{2\epsilon}dW_{t}$
$\textrm{and}\ \ \ \ d{Y}_{t}$ $=$ $\left(-\nabla_{y}c+\mathbb{E}_{p_{t}}[\nabla_{y}c|Y_{t}]+\epsilon\nabla_{y}\log f_{\nu}(Y_{t})\right)dt+\sqrt{2\epsilon}dB_{t}\ ,$ (14)
where $\epsilon>0$ is the regularization parameter; $W_{t},B_{t}$ are independent Wiener processes with values in $\mathbb{R}^{n}$; and $f_{\mu,\nu}$ are densities of the measures $\mu,\nu$. Assuming log-concave densities, Conforti et al [23] have shown that the measure generated by the Langevin dynamics (2.1) converges to the solution of (1), once the latter is regularized by entropy with parameter $\epsilon$. Despite the theoretical grounding of SDE (2.1), we have three main motivations behind focusing on OCD (2.1), instead:
  1. (1)
    While the SDE system gives us an approximation of the optimal joint density, it would not give us a map between $X_{t}$ and $Y_{t}$, due to the presence of the Brownian motion. Therefore, for applications where the random map is of interest, the SDE may not be employed; at least without further treatments.
  2. (2)
    The densities of the measures may not be known in many relevant applications. We are typically left only with samples of $\mu$ and $\nu$. Therefore the evaluation of the SDE dynamics requires further regularizations.
  3. (3)
    From the numerical aspect, computing the grad-log terms of the density, even if the latter is given, can be significantly costly if $n$ is large in $\mathbb{R}^{n}$.
The open question remains under which technical assumptions the case of $\epsilon=0$, which is OCD (2.1), would converge to the solution of the Monge-Kantorovich problem. While we do not directly address this problem and despite this knowledge gap, we present new results on both theoretical and numerical aspects of OCD, presenting its basic features along with computational relevance for practical optimal transport problems.
2.2. General Structural Properties
The OCD dynamics (2.1) leads to a certain favourable theoretical features. These are rooted in the fact that the conditional expectation is a projection operator in $L^{2}$ with self-adjoint and contraction characteristics. We leverage these characteristics to derive theoretical properties of the OCD. While the detailed justifications are deferred to Supplementary Information, we provide an overview of the main theoretical results below, where the proofs are provided in Supplementary Information.
  1. (1)
    Marginal Preservation. Since the velocities $v_{1,2}^{\textrm{ocd}}$, introduced in Eq. (2.1), are orthogonal with respect to the functions of $X_{t}$ and $Y_{t}$, respectively, a direct computation shows that the distributions of $X_{t}$ and $Y_{t}$ remain unchanged. More precisely we have the following result.
    Proposition 2.1.
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system
    $\dot{X}_{t}=v_{1}\ \ \ \textrm{and}\ \ \ \dot{Y}_{t}=v_{2}$ (15)
    with velocities $v=(v_{1},v_{2})\in\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$, and the initial condition $(X_{0},Y_{0})\sim\pi$, where $\pi\in\Pi(\mu,\nu)$ and $p_{t}$ is their joint law. Therefore $p_{t}\in\Pi(\mu,\nu)$.
  2. (2)
    Descent in Cost. As a result of the conditional expectation contraction property, the OCD leads to a monotone decay in the cost. More specifically, we have the following result.
    Proposition 2.2.
    Consider $(X_{t},Y_{t})$ to be the solution of OCD (2.1) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. We have
    $\frac{d}{dt}\mathbb{E}[c(X_{t},Y_{t})]\leq 0\ .$ (16)
    In other words, the OCD (2.1) monotonically decreases the cost $\mathbb{E}[c(X_{t},Y_{t})]$ until the latter reaches its stationary value. This property is crucial to show that the sub-optimal fixed points of the dynamics are unstable (as discussed in the following).
  3. (3)
    Instability of Sub-Optimal Couplings. The monotonic decay in the cost allows us to investigate stability of sub-optimal couplings. In particular, we have that the solution $(X_{t},Y_{t})$ of OCD with a measure from the sub-optimal coupling set
    $\Pi^{*}(\mu,\nu)\mathrel{\mathop{:}}=\left\{\pi\in\Pi(\mu,\nu)\Big{|}\int_{\mathbb{R}^{2n}}c(x,y)\pi(dx,dy)>\int_{\mathbb{R}^{2n}}c(x,y)\pi^{\textrm{opt}}(dx,dy)\right\}$ (17)
    is unstable, in the sense described below.
    Proposition 2.3.
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (2.1) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. Suppose that the joint measure of $(X_{t},Y_{t})$ fulfills
    $p_{T}\in\Pi^{*}(\mu,\nu)$ (18)
    at some $T>0$. The coupling $p_{T}$ is unstable in the sense that if the law is perturbed by
    $p_{T}^{\epsilon}$ $=$ $(1-\epsilon)p_{T}+\epsilon\pi^{\textrm{opt}}_{\mu,\nu},$ (19)
    where $10<\epsilon<$ 1, the solution $(X_{t},Y_{t})$ would not have the measure $p_{T}$ for any time later.
  4. (4)
    McKean-Vlasov System. The OCD (2.1) can be equivalently interpreted as a probability flow in the sense of Eq. (5). Due to the absence of Brownian motion, this probability flow is in the form of the Vlasov equation. The latter arises in the mean field limit of interacting particles, e.g. see [33, 49]. By exploiting the projection operator $\mathsf{S}$, the corresponding probability density evolution takes the concise form
    $\partial_{t}\rho_{t}+\nabla_{x}\cdot\left(\rho_{t}\mathsf{S}_{p_{t},x}[\nabla_{x}c]\right)+\nabla_{y}\cdot\left(\rho_{t}\mathsf{S}_{p_{t},y}[\nabla_{y}c]\right)$ $=$ $0\ ,$ (20)
    where
    $\mathsf{S}_{p_{t},x}[\nabla_{x}c]$ $=$ $\nabla_{x}c(x,y)-\mathbb{E}_{(X_{t},Y_{t})\sim p_{t}}[\nabla_{x}c(X_{t},Y_{t})|X_{t}=x]$
    $\textrm{and}\ \ \ \mathsf{S}_{p_{t},y}[\nabla_{y}c]$ $=$ $\nabla_{y}c(x,y)-\mathbb{E}_{(X_{t},Y_{t})\sim p_{t}}[\nabla_{y}c(X_{t},Y_{t})|Y_{t}=y]\ .$ (21)
    Theorem 2.4.
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (2.1) with the initial condition $(X_{0},Y_{0})\sim\pi$, from an arbitrary coupling $\pi\in\Pi(\mu,\nu)$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. Suppose $p_{t}$ is the joint law of $(X_{t},Y_{t})$ with the density $\rho_{t}$. Therefore $\rho_{t}$ is weak solution of the Vlasov-type equation Eq. (20).
    We expect that the Vlasov equation (20) admits a variational formulation in terms of Otto calculus [35, 3, 57, 9], if appropriate set of couplings is considered (those that can be generated by $\textrm{Tan}_{p_{t},2}\Pi(\mu,\nu)$). However, we leave this interesting question to a separate work. Instead, we will present a variational formulation of OCD in the probability space for the $L^{2}$ Monge-Kantorovich cost, in the next section.
2.3. Key Features in $L^{2}$ Monge-Kantorovich
So far, we discussed that OCD (2.1) brings a given joint distribution closer to the solution of the Monge-Kantorovich problem. In other words, the dynamics decreases the cost $\mathbb{E}[c(X_{t},Y_{t})]$. In this section, we present a finer result on the decay of the quadratic cost $c(x,y)=\lVert x-y\rVert_{2}^{2}$, resulted from OCD, i.e. for the ODE system
$\dot{X}_{t}=Y_{t}-\mathbb{E}_{p_{t}}[Y_{t}|X_{t}]\ \ \ \textrm{and}\ \ \ \dot{Y}_{t}=X_{t}-\mathbb{E}_{p_{t}}[X_{t}|Y_{t}]\ .$ (22)
In this scenario, the optimal transport problem is equivalent to maximizing correlation between $X_{t}$ and $Y_{t}$, see [52, 54]. Focusing on (22), we show the following results.
  1. (1)
    Symmetric Positive Semi-Definite Correlation. Consider $S^{n}_{+}$ to be the set of symmetric positive semi-definite matrices
[…]
    Let us define the cross-correlation matrix
[…]
    with $X_{t}^{\prime}\mathrel{\mathop{:}}=X_{t}-\mathbb{E}[X_{t}]$ and $Y_{t}^{\prime}\mathrel{\mathop{:}}=Y_{t}-\mathbb{E}[Y_{t}]$. We have the following result.
    Proposition 2.5.
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (2.1) with the initial condition $(X_{0},Y_{0})\sim\mu\otimes\nu$, where $\mu,\nu\in\mathcal{P}_{r}(\mathbb{R}^{n})$. Therefore
    $J_{t}\in S^{n}_{+}\ .$ (25)
    In other words the $L^{2}$-OCD (22) ensures that the cross-correlation created between $X_{t}$ and $Y_{t}$ remains in the space of symmetric positive semi-definite matrices.
  2. (2)
    Sharp Descent and Variational Formulation. The $L^{2}$-OCD (22) gives rise to a sharp decay in the cost $\mathbb{E}[c(X_{t},Y_{t})]$. More precisely, if we restrict the set of velocities to the tangent space $(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}$, the OCD gives fastest decay in the cost $\mathbb{E}[c(X_{t},Y_{t})]$.
    Theorem 2.6.
    Consider $(X_{t},Y_{t})$ to be the solution of the ODE system (15) and the initial condition $(X_{0},Y_{0})\in\pi$, where $\pi\in\Pi(\mu,\nu)$. Let
[…]
    be the $L^{2}$ transport cost and
[…]
    its decay rate penalized by the $L^{2}$ norm of the velocities. Define optimal velocities
[…]
    Therefore
    $v_{p_{t},1}^{opt}(X_{t},Y_{t})$ $=$ $Y_{t}-\mathbb{E}_{p_{t}}[Y_{t}|X_{t}]$ (29)
    $\textrm{and}\ \ \ \ v_{p_{t},2}^{opt}(X_{t},Y_{t})$ $=$ $X_{t}-\mathbb{E}_{p_{t}}[X_{t}|Y_{t}]\ .$ (30)
    This result can be employed as variational formulation of $L^{2}$-OCD. The velocities $v_{1,2}^{\textrm{ocd}}$ admit the variational formulation
    $v_{1,2}^{\textrm{ocd}}$ $=$ $\operatorname*{arg\,min}_{v\in(L^{2}(\mu,\mathbb{R}^{n})\otimes L^{2}(\nu,\mathbb{R}^{n}))^{\perp}}\left[\frac{d}{dt}\int_{\mathbb{R}^{2n}}\lVert x-y\rVert_{2}^{2}p_{t}(dx,dy)+\int_{\mathbb{R}^{2n}}\lVert v\rVert_{2}^{2}p_{t}(dx,dy)\right]\ .$ (31)
3. Special Examples for $L^{2}$-Monge-Kantorovich
We present two scenarios for which the solution of OCD can be analyzed with respect to the analytical solution of $L^{2}$-Monge-Kantorovich problem.
3.1. Linear Maps
After initializing $(X_{0},Y_{0})\sim\mu\otimes\nu$, suppose $L^{2}$-OCD (22) converges to a stationary solution $(X_{s},Y_{s})$ which is represented by an affine map
$Y_{s}=T(X_{s})=\Sigma X_{s}+m\ ,$ (32)
with $\Sigma\in\mathbb{R}^{n\times n}$ and $m\in\mathbb{R}^{n}$. We argue that the map $T()$ is the Monge (optimal) map from $\mu$ to $\nu$. Due to Proposition 2.5, $\Sigma\in S_{+}^{n}$, i.e. it has to be symmetric positive semi-definite. Therefore the linear map can be cast as the gradient
$T(X_{s})$ $=$ $\nabla_{x}\phi(X_{s})$ (33)
of a convex
[…]
It follows from Brenier’s theorem [16, 2, 56] that a map is solution of $L^{2}$-Monge-Kantorovich problem if (and only if) it is gradient of a convex. Therefore $T$ is the Monge map.
3.2. Elliptic Distributions
A special case is when both $\mu$ and $\nu$ are Gaussian measures. Let $\mu$ and $\nu$ be centered Gaussian measures with covariance matrices $\Sigma_{\mu}$ and $\Sigma_{\nu}$. The dynamics of correlation matrix $J_{t}=\mathbb{E}[X_{t}\otimes Y_{t}]$, following $L^{2}$-OCD (2.1), reads
$\partial_{t}J_{t}$ $=$ $\Sigma_{\mu}+\Sigma_{\nu}-J_{t}^{T}\Sigma_{\mu}^{-1}J_{t}-J_{t}\Sigma_{\nu}^{-1}J_{t}^{T}\ ,$ (35)
due to the identities
$\mathbb{E}_{p_{t}}[Y_{t}|X_{t}]=J_{t}^{T}\Sigma_{\mu}^{-1}X_{t}\ \ \ \textrm{and}\ \ \ \mathbb{E}_{p_{t}}[X_{t}|Y_{t}]=J_{t}\Sigma_{\nu}^{-1}Y_{t}\ ,$ (36)
which results from $p_{t}$ being Gaussian, see [23]. The matrix-Riccati equation (35) admits many stationary solutions, since any bijection between $\mu$ and $\nu$ is a stationary solution of (2.1), as discussed in [23]. However, as we justified in Proposition 2.2, starting with independent samples, i.e. $\mathbb{E}[X_{0}Y_{0}]=0$, $J_{t}$ remains symmetric positive semi-definite and its trace increases monotonically. This restricts the stationary solution to the one associated with the optimal transport. More precisely, we have
$\dot{X}_{t}=Y_{t}-J_{t}^{T}\Sigma_{\mu}^{-1}X_{t}\ \ \ \textrm{and}\ \ \ \dot{Y}_{t}=X_{t}-J_{t}\Sigma_{\nu}^{-1}Y_{t}$ (37)
with the fixed point
$Y_{s}=J_{s}^{T}\Sigma_{\mu}^{-1}X_{s}\ \ \ \textrm{and}\ \ \ X_{s}=J_{s}\Sigma_{\nu}^{-1}Y_{s}\ .$ (38)
Given that $J_{s}$ is symmetric positive semi-definite (Proposition 2.2), each of $X_{s}$ or $Y_{s}$ can be expressed as gradient of a convex function of the other (similar to Eq. (34)). For the one-dimensional case, Eq. (35) takes a simpler form. Let $n=1$ and suppose the variances are $\sigma^{2}_{\mu}$ and $\sigma^{2}_{\nu}$. Define the correlation coefficient $\kappa_{t}\mathrel{\mathop{:}}=J_{t}/(\sigma_{\mu}\sigma_{\nu})$. Therefore, the Riccati equation (35) reduces to
$\partial_{t}\kappa_{t}$ $=$ $\left(\frac{\sigma_{\mu}^{2}+\sigma_{\nu}^{2}}{\sigma_{\mu}\sigma_{\nu}}\right)\left(1-\kappa_{t}^{2}\right)\ .$ (39)
Starting from independent samples, this nonlinear ODE gives a monotone increase of $\kappa_{t}$ from zero to the stationary solution $\kappa_{t}\to 1$ as $t\to\infty$, which is consistent with the solution of the optimal transport problem (note that here the Monge map is given by $\hat{T}(X)=\sigma_{\nu}/\sigma_{\mu}X$ leading to unity correlation coefficient).
4. Numerical Algorithm through Opinion Dynamics
Opinion dynamics at its core tries to answer how different agents shape their opinions in contact with each-other [37]. While seemingly unrelated to the Monge-Kantorovich problem, in this section, we leverage analogy between opinion dynamics and OCD to provide a novel, yet simple, numerical algorithm for the latter. The main idea is to first, split the data points into a certain number of clusters. Next, the conditional expectation is estimated in each cluster using an appropriate hypothesis class.
4.1. Analogy with Opinion Dynamics
Consider a population of agents indexed by $i\in\{1,...,N\}$. Suppose each person $i$ has certain two-dimensional position $P_{i}(t)\in\mathbb{R}^{2}$ as their opinion in time $t$. In the standard form of consensus dynamics, e.g. see [44], these positions relax towards a global consensus via
$\dot{P}_{i}$ $=$ $\sum_{i\neq j}a_{ij}(P_{i}(t)-P_{j}(t))\ ,$ (40)
where $a_{ij}\geq 0$ captures the influence of other agents positions. We can decompose it as $a_{ij}=\phi(r_{ij})\alpha_{ij}$ where $r_{ij}=\lVert P_{i}-P_{j}\rVert_{2}^{2}$ is the Euclidean distance between the positions of two persons $i$ and $j$, and $\alpha$ a constant matrix. Based on $\alpha$, we can intuitively consider two idealized regimes:
  1. (1)
    The positions are completely independent from each-other and $\alpha$ becomes the identity matrix.
  2. (2)
    The positions are perfectly tied to each-other and $\alpha$ becomes the exchange matrix $J$.
The first one degenerates the system into independent positions. More interestingly, let us focus on the latter and denote $P=[P_{X},P_{Y}]^{T}$, therefore the dynamics takes the form
$\dot{P}_{X,i}=\sum_{i\neq j}\phi(r_{ij})(P_{Y,i}(t)-P_{Y,j}(t))\ \ \textrm{and}\ \ \dot{P}_{Y,i}=\sum_{i\neq j}\phi(r_{ij})(P_{X,i}(t)-P_{X,j}(t)).$ (41)
If we assume $\phi(\ )$ is simply a step function i.e. it is one if $r\leq\epsilon$ and zero otherwise (for $\epsilon>0$), we can further simplify the dynamics to
$\dot{P}_{X,i}=\frac{1}{N_{\mathrm{bar}{I}^{X_{i},\epsilon}}}\sum_{j\in\mathrm{bar}{I}^{X_{i},\epsilon}}(P_{Y,i}(t)-P_{Y,j}(t))\ \ \textrm{and}\ \ \dot{P}_{Y,i}=\frac{1}{N_{\mathrm{bar}{I}^{Y_{i},\epsilon}}}\sum_{j\in\mathrm{bar}{I}^{Y_{i},\epsilon}}(P_{X,i}(t)-P_{X,j}(t))\ ,$ (42)
where $\mathrm{bar}{I}^{X_{i},\epsilon}$ and $\mathrm{bar}{I}^{Y_{i},\epsilon}$ are the set of all agents with $\lVert P_{X,i}-P_{X}\rVert_{2}^{2}\leq\epsilon^{2}$ and $\lVert P_{Y,i}-P_{Y}\rVert_{2}^{2}\leq\epsilon^{2}$, respectively, with the corresponding number of elements $N_{\mathrm{bar}{I}^{X_{i},\epsilon}}$ and $N_{\mathrm{bar}{I}^{Y_{i},\epsilon}}$. In the limit of $N\to\infty$ and $\epsilon\to 0$, it is suggestive to express the above dynamics in the statistical form. Treating $P_{X,i}(t)$ and $P_{Y,i}(t)$ as samples of random variables $X_{t}\in\mathcal{H}$ and $Y_{t}\in\mathcal{H}$, respectively, the dynamics takes the form
$\dot{X}_{t}=Y_{t}-\mathbb{E}[Y_{t}|X_{t}]\ \ \ \textrm{and}\ \ \ \dot{Y}_{t}=X_{t}-\mathbb{E}[X_{t}|Y_{t}]\ ,$ (43)
which is OCD (2.1) for $L^{2}$-cost. From technical point of view, the discrete version is justified due to the approximation of conditional expectation by a discrete kernel [46]. Through the lens of opinion dynamics, the OCD describes a dynamics which is local in a sense that each agent is only influenced by agents in its vicinity, yet as discussed before, it is closely linked to the global optimization problem (1).
4.2. Non-parametric Monte-Carlo Algorithm
The analogy with opinion dynamics hints a numerical recipe for OCD. Conceptually, the parameter $\epsilon$ gives rise to formation of clusters in the cloud of $(X_{t},Y_{t})$ points. We consider two numerical schemes:
  1. (1)
    OCD-piecewise constant. In each cluster, $X_{t}$ and $Y_{t}$ relax towards their locally averaged values, i.e. the conditional expectation is estimated to be constant in the cluster.
  2. (2)
    OCD-piecewise linear. In each cluster, the conditional expectation is estimated using linear regression, i.e. the joint distribution of each cluster is estimated by a Gaussian.
Starting with $N_{p}$ i.i.d. samples $\hat{X}_{i,t_{0}}\sim\mu$ and $\hat{Y}_{i,t_{0}}\sim\nu$, we update the state of each sample according to discretized form of OCD. It is important to note that here we are focusing on an approximate solution of the optimization problem (1) over atom-less distributions $\mu$ and $\nu$. An alternative interpretation, where the distributions are assumed to be sum of diracs (assignment problem) is not pursued here (see [62]). Therefore throughout what follows, we assume that there exists an underlying continuous map between the two measure spaces. Three hyper-parameters control the discretization error: number of samples $N_{p}$, time step size $\Delta t$, and cluster cut-off $\epsilon$. Once these values are fixed, relying on Euler’s scheme (see Runge-Kutta’s scheme in Sec.7.2 of Supplementary Information), each sample is updated according to
$\hat{X}^{n+1}_{i}$ $=$ $\hat{X}^{n}_{i}+\nabla_{x}c(\hat{X}_{i}^{n},\hat{Y}_{i}^{n})\Delta t-\hat{\mathcal{K}}^{n}_{X}\Delta t$ (44)
$\textrm{and}\ \ \ \hat{Y}^{n+1}_{i}$ $=$ $\hat{Y}^{n}_{i}+\nabla_{y}c(\hat{X}_{i}^{n},\hat{Y}_{i}^{n})\Delta t-\hat{\mathcal{K}}^{n}_{Y}\Delta t\ ,$ (45)
where superscript $n$ and $n+1$ denote the approximations at $t^{n}$ and $t^{n}+\Delta t$, respectively. The functionals $\hat{\mathcal{K}}^{n}_{X,Y}(.)$ estimate conditional expectations with respect to $X$ and $Y$, respectively, and they are found depending on the following estimators.
  1. (1)
    OCD-piecewise constant. The estimator is constant in each cluster (equivalent to using square kernel in the estimation of the conditional expectation). The expressions for $\hat{\mathcal{K}}^{n}_{X,Y}(.)$ are given in Eqs. (76)-(77) in Supplementary Information.
  2. (2)
    OCD-piecewise linear. The estimator follows linear regression (accounts for using Gaussian kernel in the estimation of the conditional expectation). In other words here we look for $L^{2}$ projection of $\nabla_{x,y}c$ onto the span of linear functions [36]. The final expressions are given in Eqs. (85)-(86) in Supplementary Information.
The overview of the algorithm is given in Algorithm 1 in Supplementary Information. The convergence of the algorithm, in general, is controlled by $N_{p},\Delta t,\epsilon$. Assuming that the OCD initialized by samples of $\mu,\nu$ is well-posed, it is reasonable to consider fixed-order convergence with respect to $\Delta t$ (e.g. first-order for Euler time integration, see [50]). Also the error due to finite $N_{p}$ is expected to behave similar to conventional Monte-Carlo schemes [38]. However, the error dependency on the cut-off value $\epsilon$ is more delicate. Two asymptotic limits of $\epsilon\to 0$ and $\epsilon\to\infty$ can be intuitively understood. The limit $\epsilon\to 0$ accounts for no interaction between agents, i.e. each point forms its own cluster. Therefore the points remain frozen in time. On the other hand, $\epsilon\to\infty$ gives rise to perfect interaction among all agents, forming one giant cluster. In this case, the dynamics converges to the average value among all sample points, for the 1st order approximation, and towards linear regression, for the 2nd order case. Obviously, the optimal value of $\epsilon$ needs to be tuned in-between, and is expected to depend on $N_{p}$. The transition between frozen dynamics in $\epsilon\to 0$ and homogenization $\epsilon\to\infty$ suggests a phase transition type behaviour, resulting from $\epsilon$. The transition point can serve as a proxy for the optimal cut-off value, as discussed in Sec. 7.3 in Supplementary Information.
Once the pairing between $\hat{X}_{i}$ and $\hat{Y}_{i}$ reaches its stationary state (according to a stopping criterion), the sampled points $(\hat{X}_{i}^{s},\hat{Y}_{i}^{s})_{i\in\{1,...,N_{p}\}}$ can be employed to estimate the Monge map $\hat{T}()$, using a suitable hypothesis class. We train a Neural Network (NN) with a few layers and $\tanh(.)$ as the activation function to approximate the map and denote it by $M_{X\rightarrow Y}$. As the loss function, we simply consider the expectation of $L^{2}$ point-wise error between NN’s output $M_{X\rightarrow Y}(\hat{X}^{*})$ and $\hat{Y}^{*}$, i.e. $\mathrm{Loss}=\sum_{i=1}^{N_{p}}\lVert M_{X\rightarrow Y}(\hat{X}_{i}^{*})-\hat{Y}_{i}^{*}\rVert_{2}^{2}/N_{p}$.
4.3. Computational Complexity
For a given time step (iteration), the main computational task of the proposed OCD algorithm is evaluation of the conditional expectation, whose estimation relies on forming clusters by finding neighbours of each point in a given distribution. For the latter, though non-exclusively, the Ball Tree method [47] scales efficiently i.e. $\mathcal{O}(N_{p}\log N_{p})$ with number of particles $N_{p}$. Furthermore, the cost of OCD scales linearly with $n=\dim(X)=\dim(Y)$, since the number of ODEs is $2n$ and number of operations in order to compute the distance between each two particles is $\mathcal{O}(n)$. In terms of memory consumption, since OCD, unlike linear programming, does not require computation of the distance matrix, its memory requirement is $\mathcal{O}(N_{p})$. In order to estimate the total computational cost of OCD, including number of iterations required to reach stationary state, we conducted numerical tests as shown in Fig. 2, where OCD runs between two normal distributions. The flops and storage requirements confirm our analysis on complexity and memory consumption. However, it is evident that OCD requires significantly more operations compared to the reference Earth Mover’s Distance (EMD) method [29], if relatively small number of sample points (e.g. 100) is considered. This is mainly due to the fact that while the number of operations scales efficiently, i.e. with $\mathcal{O}(N_{p}\log N_{p})$, the pre-factor which depends on the number of iterations required to reach stationary condition becomes dominant, if $N_{p}$ is relatively small.
Refer to caption Refer to caption
Refer to caption Refer to caption
(a) $\dim(X)=3$ (b) $N_{p}=800$
Figure 2. Execution time (top) and memory consumption (down) of OCD-piecewise linear algorithm with Runge Kuta 4th order solver with $\Delta t=0.1$ against Earth Mover’s Distance (EMD) method [29] in learning the map between samples of $\mu=N(0,I)$ and $\nu=N(1,I)$ in (a) $\dim(X)=3$ for a range of $N_{p}$ and (b) $\dim(X)=1,...,8$ with $N_{p}=800$. Here we scale $\epsilon$ with a rule of thumb $\varepsilon=3d{N_{p}}^{-1/4}/4$.
5. Representative Results for $L^{2}$-Monge-Kantorovich
In this section, we illustrate the accuracy and cost of the proposed OCD and highlight its capability in estimating the optimal map, along with Wasserstein distance, between two marginals in a few examples. Unless mentioned otherwise, by OCD we are referring to OCD-piecewise linear algorithm equipped with Runge-Kutta 4th order ODE solver. We refer the reader to Sec.7.4 in Supplementary Information for details on the choice of hyper-parameters such as $\epsilon$ and $\Delta t$, and comparison with benchmarks. All the performance studies are done on a single core and thread of Intel Core i7-8550U CPU with 1.80 GHz frequency.
5.1. Clustering and Prototypical Examples
Figure 3. Evolution of point particles under the coupling created by OCD.
Figure 4. Evolution of cross-correlation matrix, $J_{t}$, in the cone of symmetric positive semi-definite matrices.
Figure 5. Recovery of nonlinear Monge map between (a) normal distribution and (b) its push-forward by softmax map $T_{i}=\nabla_{x_{i}}\log\left(\exp(x_{1})+\exp(x_{2})\right)$; (c) Starting from independent samples, the pair $(X_{t},Y_{t})$ is evolved by OCD and $\mathbb{E}\lVert X_{t}-Y_{t}\rVert_{2}^{2}$ converges towards the Wasserstein distance $d^{2}$; (d) First component of the map generated by OCD; (e) First component of the exact Monge map.
Fist, we present and discuss three toy scenarios.
  1. (1)
    Particle Clustering. Starting from independent samples of two one-dimensional standard normal distributions, we track the particles evolved by $L^{2}$-OCD. The dynamics is illustrated in Fig. 3, where the correlation between the data points monotonically increases in time. The Monge map, given by identity, emerges from particle clusters as time progresses.
  2. (2)
    Symmetric Positive Semi-Definite Cone. As a by-product of OCD, we justified in Proposition 2.5 that the dynamics ensures the cross-correlation matrix remains in the cone of symmetric positive semi-definite matrices. Here we verify this finding by numerical simulation, where the initial condition is product measure of two centered Gaussians with covariance matrices $\Sigma_{\mu}$ and $\Sigma_{\nu}$. As shown in Fig. 4, OCD evolves the cross-correlation inside the cone and reaches the maximum on the curve connecting $\Sigma_{\mu}$ and $\Sigma_{\nu}$. Note that there exist interesting results on the geometry of the cross-correlation of Gaussians induced by the $L^{2}$-optimal transport and its connection to geometric average of $\Sigma_{\mu}$ and $\Sigma_{\nu}$, which we leave out of discussion for brevity, but refer the reader to [11].
  3. (3)
    Nonlinear Monge Map. To demonstrate relevance of OCD in recovering nonlinear multi-dimensional maps, we consider a synthetic setting where the target distribution is chosen such that the Monge map is known analytically. Consider the standard normal distribution over $\mathbb{R}^{2}$, denoted by $\mu$. Let the target distribution be the push-forward by softmax map $\nu=T\#\mu$, with $T=\nabla_{x}\log(\exp(x_{1})+\exp(x_{2}))$. Since $T$ is gradient of a convex, the Monge map between $\mu$ and $\nu$ is simply $T$. Depicted in Fig. 5, starting from independent $10^{6}$ samples of $\mu$ and $\nu$, particles pushed by OCD with $\epsilon=10^{-3}$ and $\Delta t=10^{-2}$ successfully recover the Monge map and accurately estimate 2-Wasserstein distance, as $t$ progresses.
5.2. Distribution Learning and Sampling
[…]
For testing, we generate $10^{6}$ normally distributed samples, i.e. $X^{\text{test}}\sim\mathcal{N}(0,I_{n\times n})$ and feed them to the NN to find $M_{X\rightarrow Y}(X^{\text{test}})$. As shown in Fig. 6, OCD is capable of recovering the target density with a reasonable accuracy. However, we note that the estimated densities by OCD are slightly smoothed out compared to the ground-truth. This error could be related to the current approach of estimating the conditional expectation, which resembles Kernel Density Estimation with known smoothing effects. In Sec. 7.4 of Supplementary Information, we provide a study of hyper-parameters and comparison with Adaptive Transport Maps (ATM) [8] for each test case. We point out that once the map is learned, computing the new samples is straight-forward, given that the network only performs simple matrix-vector multiplication on the samples of the standard normal distribution. Hence the proposed workflow can efficiently serve as a sampling approach for complex distributions (such as Swiss roll), with ramifications for sampling algorithms based on measure transport [60, 48].
True
Refer to caption Refer to caption Refer to caption
OCD
Refer to caption Refer to caption Refer to caption
Normal Banana Funnel Swiss roll
Figure 6. Estimating density of Banana, Funnel, and Swiss roll distributions using OCD. Here, first we find optimal pairs among $10^{4}$ samples of normal and target density using OCD. We train NN on the sorted data and generate new samples of the target density using the learned map and $10^{6}$ normally distributed particles. The true marginals are depicted on top and histogram of estimated marginals on bottom.
5.3. Classification in a Data Set
Since the Wassetstein distance provides a metric in the space of distributions, it can be leveraged to label data-sets in the context of machine-learning. For example, it allows for grouping similar data points and finding out their labels. To showcase the relevance of OCD in such applications, we compute the 2-Wasserstein distance in the Japanese Female Facial Expression (JAFFE) data-set [42]. As shown in Fig. 7, the proposed OCD (with $L^{2}$-cost) can find similarities between images. In particular, the distance between different images corresponding to the same model takes minimal values, on the diagonal elements. It is encouraging to see that given negligible memory requirement of OCD, we could carry out the computations on a local machine in a few seconds with fraction of memory requirement compared to EMD: for two $128\times 128$ pixel pictures, EMD consumes $6.5$ GB while OCD only needs $181$ MB of memory.
Refer to caption Refer to caption
Distribution of Wasserstein distance Example of found similar pictures
Figure 7. Distribution of the Wasserstein distance (left) and example of found similar pictures with minimum distance in each row (right) among $64\times 64$ pixels pictures in the Japanese Female Facial Expression (JAFFE) dataset using OCD with $\Delta t=0.1$ and $\epsilon=\epsilon_{\textrm{max}}(N_{\textrm{clusters}}/N_{p}>0.99)$.
5.4. Color Interpolation
As the final example, we consider another application of optimal transport in interpolation between color distribution of different images. Given an image with dimension $(N_{x},N_{y},3)$, the pixel colors can be considered as the random variable, leading to $N_{x}\times N_{y}$ samples for $X$ with $\dim(X)=3$, indicating red/green/blue channels. We take a low resolution picture of the ocean taken in day and sunset time for training, as depicted in Fig. 8 (see examples of POT library [29] for details).
Refer to caption Refer to caption
$X_{0}$ $Y_{0}$
Figure 8. Low resolution ($84\times 125$ pixels) pictures of ocean in day and sunset time used for training.
Given $N_{p}=10^{\prime}500$ training samples of these two pictures, we learn the map $M_{X\rightarrow Y}$ using paired samples found from OCD (with $L^{2}$-cost) and training NN similar to the one described in 5.2. We employ $\Delta t=0.1$ and $\epsilon=\epsilon_{\textrm{max}}(N_{\textrm{clusters}}/N_{p}>0.9)$. We repeat the same procedure for EMD to benchmark, and found that while EMD takes about $2.6$ GB of memory, OCD only consumes $115$ MB. Once the NN map is learned, we test the map by feeding the NN with the full resolution of day picture with $670\times 1000$ pixels, where we see a good agreement with the EMD solution. We further test the learned OCD map on two new testing picture with $548\times 1024$ and $1200\times 1920$ pixels. As shown in Fig. 9, the trained OCD map is capable of transforming the testing picture into sunset.
OCD
EMD
Test #1: OCD
Test #2: OCD
Figure 9. Interpolating day and sunset pictures of ocean. The interpolation is done using the learned map $M_{X\rightarrow Y}(.)$ via $(1-\alpha)X+\alpha M_{X\rightarrow Y}(X)$. The bottom two rows show the transport of colors for two test picture.
6. Discussion
Conceptually, what makes the OCD approach particularly appealing is that it is at cross-road of different topics:
  1. (1)
    The original problem of Monge-Kantorovich optimization, constrained on a set of couplings, is reduced to a dynamics whose evolution is controlled by the conditional expectation. The latter is solution of the regression problem, and thereby, the OCD provides a direct link between Bayesian regression and the optimal transportation.
  2. (2)
    Using basic interaction rules, the dynamics gives a particle-type relaxation of independent samples, seeking their optimal coupling. The notion of clustering, which arises from this coupling, resonates with models of opinion dynamics and offers a method for obtaining sparse representations of complex, high-dimensional datasets.
  3. (3)
    The OCD framework has potential applications as the foundation for generative models. When combined with supervised learning, the learned maps could be employed to generate new datasets.
  4. (4)
    Finally, the dynamics of OCD can be interpreted as a form of projected gradient descent, providing a novel approach through which optimization techniques constrained on a set of distributional coupling can be constructed.
From numerical point of view, the model has favorable scaling behaviour. In the most basic form, it scales linearly with number of data points as well as dimension of distributional support, making OCD competitive in comparison to the state-of-the-art random map algorithms as well as direct linear programming approach. This powerful computational efficiency enables us to treat distributional learning problems over more complex and higher dimensional settings.
Limitations. The current treatment of conditional expectation exhibits diffusive behaviour which can deteriorate marginal distributions with sharp features. More elaborated approaches, e.g. deployment of Gaussian Processes or Neural Nets as hypothesis class of the conditional expectation, may alleviate this issue, at least to some degree, though likely at the cost of increased computational complexity. In addition, moment constraints of the marginal distributions can be leveraged to enforce the marginal densities of OCD to remain close to the initial condition, and hence reducing the numerical dispersion. Another issue that requires further investigation is the optimal choice of $\epsilon$. Although we carried out extensive numerical tests, see Sec.  7.3 in Supplementary Information, we believe computing the conditional expectation efficiently is a separate task by itself. These trade-offs between accuracy and efficiency are interesting areas for future research on OCD. From a theoretical standpoint, our understanding of the underlying dynamics remains incomplete. While we provided some important theoretical results on OCD, questions such as the well-posedness of OCD and its convergence towards optimal solution are still unresolved and require further exploration.
Broader Context. The presented dynamics, along with devised non-parametric basis-free algorithm, gives a means to pair up distinct data-sets, through a dynamics whose increment is a functional of the distribution. This is a prototype of McKean-Vlasov processes, which offer further generalizability, compared to the conventional regression, due its non-local dependency on the distribution. We believe this development holds implications which could go beyond the context of optimal transport, specifically given recent attention towards McKean-Vlasov universal approximators emerging in data-driven trainable models [64, 6].